Find, showing all working, a formula for the n-th term tn of the sequence (tn) defined by

t1 = 5; tn = -7tn-1 /3, n >= 2.


1
Expert's answer
2021-10-06T14:07:09-0400
"t_1=5, t_n=-\dfrac{7t_{n-1}}{3}"

"\dfrac{t_n}{t_{n-1}}=\dfrac{-\dfrac{7t_{n-1}}{3}}{t_{n-1}}=-\dfrac{7}{3}"

We have a geometric progression "(t_n)"


"t_1=5, q=-\dfrac{7}{3}"

Then


"t_n=5(-\dfrac{7}{3})^{n-1}, n\geq1"